# Logarithmic regret in the ergodic Avellaneda–Stoikov market making model

### v1

@@ 2.2. Learning and regret
There exist a constant $C>0$ such that
$0\leq\gamma(\kappa^{\star};\kappa^{\star})-\gamma(\kappa;\kappa^{\star})\leq C\left|\kappa-\kappa^{\star}\right|^{2},\,\forall\kappa\in[\underline{K},\bar{K}]\,.$
So far we have performed the ergodic analysis for the market making model with parameter $\kappa$ misspecified. Another key step towards quantifying the performance gap, see Theorem 27, is to analyse the ergodicity under the model misspecification. By simply substituting $\kappa$ for $\kappa^{\star}$ and $\delta^{*,\pm}$ for $\psi^{\kappa}$ in the state transition matrix $\bm{P}$ defined in Appendix A.10, we can conclude that there still exists a unique equilibrium, denoted by $\pi^{\psi^{\kappa};\kappa^{\star}}$, for $Q_{t}^{\psi^{\kappa};\kappa^{\star}}$ following (22). Proposition 26, whose proof is provided in Appendix A.16, gives a nice property of the equilibrium $\pi^{\psi^{\kappa};\kappa^{\star}}$.
@@
Let $\pi^{\psi^{\kappa};\kappa^{\star}}$ be the equilibrium distribution for the inventory process $Q_{t}$ following the controlled SDE (22)
$dQ_{t}^{\psi^{\kappa};\kappa^{\star}}=\big{(}\lambda^{+}e^{-\kappa^{\star}\psi^{\kappa,-}}-\lambda^{-}e^{-\kappa^{\star}\psi^{\kappa,+}}\big{)}dt+d\tilde{N}_{t}^{\psi^{\kappa,-}}-d\tilde{N}_{t}^{\psi^{\kappa,+}},\quad t\in[0,T],Q_{0}\sim\pi^{\psi^{\kappa};\kappa^{\star}},$
@@
$\begin{split}&\mathbb{E}\bigg{[}\int_{0}^{T}\lambda^{+}e^{-\kappa^{\star}\psi^{\kappa,+}(Q_{t}^{\psi^{\kappa};\kappa^{\star}})}\Big{(}\hat{v}^{\psi^{\kappa}}(Q_{t}^{\psi^{\kappa};\kappa^{\star}}-1;\kappa^{\star})-\hat{v}^{\psi^{\kappa}}(Q_{t}^{\psi^{\kappa};\kappa^{\star}};\kappa^{\star})\Big{)}\bm{1}_{Q_{t}^{\psi^{\kappa};\kappa^{\star}}>\underline{q}}\\ &+\lambda^{-}e^{-\kappa^{\star}\psi^{\kappa,-}(Q_{t}^{\psi^{\kappa};\kappa^{\star}})}\Big{(}\hat{v}^{\psi^{\kappa}}(Q_{t}^{\psi^{\kappa};\kappa^{\star}}+1;\kappa^{\star})-\hat{v}^{\psi^{\kappa}}(Q_{t}^{\psi^{\kappa};\kappa^{\star}};\kappa^{\star})\Big{)}\bm{1}_{Q_{t}^{\psi^{\kappa};\kappa^{\star}}<\bar{q}}\,\mathrm{d}t\bigg{]}=0\,,\end{split}$
where $\hat{v}^{\psi^{\kappa}}(q;\kappa^{\star})$ is defined in Proposition 21.
@@
$\mathcal{R}^{\Psi}(T)\leq C_{1}\mathbb{E}\Big{[}\int_{0}^{T}\left|\kappa_{t}-\kappa^{\star}\right|^{2}\,\mathrm{d}t\Big{]}+C_{2}\frac{e^{T\ln\alpha}-1}{\ln\alpha},$
@@
The next step towards Theorem 19 is to quantify the precise tail behaviour, also known as concentration inequality, of the regularised maximum likelihood estimator in Algorithm 1. Recall that $(\Omega^{\star},\mathcal{F}^{\star},\mathbb{P}^{\star})$ is given in Definition 14.
We start with several significant propositions to the regularised maximum likelihood estimator. The proofs of the following propositions are provided in Appendix A.18.
@@
Let $(\delta_{n})_{n=1}^{N}$ be a collection of non-negative random variables taking values in $[\underline{\delta},\bar{\delta}]\cup\{+\infty\}$. Then for any $\epsilon\geq 0$ and bounded function $f:[\underline{\delta},\bar{\delta}]\cup\{+\infty\}\rightarrow\mathbb{R}$, it holds that,
$\mathbb{P}^{\star}\left(\left|\sum_{n=1}^{N}f(\delta_{n})Y_{n}-\sum_{n=1}^{N}f(\delta_{n})e^{-\kappa^{\star}\delta_{n}}\right|\leq\left\|f\right\|_{\infty}\sqrt{2N\ln(\tfrac{2}{\epsilon})}\right)\geq 1-\epsilon.$
Remark 29.
Note that the estimator $\kappa$ generated by Algorithm 1 is bounded in the set $[\underline{K},\bar{K}]$, hence it is not hard to see that the optimal ergodic control $\delta$ takes values in $[\underline{\delta},\bar{\delta}]\cup\{+\infty\}$ with $\underline{\delta},\bar{\delta}\in\mathbb{R}$ by (18).
@@
There exist constants $c,C>0$ depending on $\bar{K},\underline{\delta}$ and $\bar{\delta}$ such that for any policy $(\delta_{n})_{n=1}^{\infty}$ taking values in $[\underline{\delta},\bar{\delta}]\cup\{+\infty\}$, it holds that for any $\epsilon>0$,
$\mathbb{P}^{\star}\left(\inf_{\kappa\geq 0}\Big{(}-\frac{\mathrm{d}^{2}}{\mathrm{d}\kappa^{2}}\tilde{\ell}_{N}(\kappa)\Big{)}\geq cN-C\sqrt{N\ln\big{(}\tfrac{2}{\epsilon}\big{)}}\right)\geq 1-\epsilon.$
@@
There exists constants $C,c\geq 0$ depending on $\kappa^{\star}$, $\delta_{0}$, and $\bar{\delta}$ such that for any policy $(\delta_{n})_{n=1}^{N}$ taking values in $[\underline{\delta},\bar{\delta}]\cup\{+\infty\}$, it holds that for any $\epsilon>0$,
$\mathbb{P}^{\star}\left(\Big{|}\frac{\mathrm{d}}{\mathrm{d}\kappa}\tilde{\ell}_{N}(\kappa^{\star})\Big{|}\leq C\sqrt{N\ln\big{(}\tfrac{2}{\epsilon}\big{)}}+c\right)\geq 1-\epsilon.$
@@
Let $\kappa_{N}>0$ be the unique solution to $\frac{\mathrm{d}}{\mathrm{d}\kappa}\tilde{\ell}_{N}(\kappa_{N})=0$. There exists constants $C,c,N_{0}\geq 0$ such that for any $\epsilon\geq 0$, if $N\geq N_{0}\ln\big{(}\tfrac{2}{\epsilon}\big{)}$, then
$\mathbb{P}^{\star}\left(|\kappa_{N}-\kappa^{\star}|\leq CN^{-1/2}\sqrt{\ln\big{(}\tfrac{2}{\epsilon}\big{)}}+cN^{-1}\right)\geq 1-2\epsilon.$
@@
Let $\kappa_{N}>0$ be the unique solution to $\frac{\mathrm{d}}{\mathrm{d}\kappa}\tilde{\ell}_{N}(\kappa_{N})=0$. Then there exists constants $C,c,N_{0}\geq 0$ such that for any $\epsilon\geq 0$,
$\mathbb{P}^{\star}\left(|\kappa_{N}-\kappa^{\star}|\leq CN^{-1/2}\sqrt{\ln\big{(}\tfrac{2N}{\epsilon}\big{)}}+cN^{-1}\quad\text{for all}\quad N\geq N_{0}\ln\big{(}\tfrac{2}{\epsilon}\big{)}\right)\geq 1-\epsilon\,.$
@@ 2.2.1. Step 3: Proof of Theorem 19
$\int_{0}^{T}|\kappa_{t}-\kappa^{\star}|^{2}\,\mathrm{d}t\leq\sum_{n=0}^{N_{T}}(\tau_{n+1}-\tau_{n})|\kappa_{n}-\kappa^{\star}|^{2}=:X_{N_{T}}\,.$
@@
The following proposition, Proposition 34, which is proved in Appendix  A.20, states that, given any $N_{T}$, i.e. the number of signals up to time $T$, the random variable $X_{N_{T}}$ is bounded by $\mathcal{O}(\ln^{2}N_{T})$ with high probability.
@@
There exist constants $C_{1},C_{2},C_{3},C_{4}>0$ such that for any $\epsilon>0$,
$\mathbb{P}^{\star}\left(X_{N_{T}}\leq C_{1}\ln^{2}N_{T}+C_{2}\ln N_{T}\ln(\tfrac{2}{\epsilon})+C_{3}\ln^{2}(\tfrac{2}{\epsilon})+C_{4}\right)\geq 1-2\epsilon.$
Let $r(N_{T}):=C_{1}\ln^{2}N_{T}+C_{2}\ln N_{T}\ln(\tfrac{2}{\epsilon})+C_{3}\ln^{2}(\tfrac{2}{\epsilon})+C_{4}$. Then, by using Proposition 34, we have
$\begin{split}\mathbb{E}[X_{N_{T}}]&=\mathbb{E}\Big{[}\mathbb{E}\big{[}X_{N_{T}}\big{|}N_{T}\big{]}\Big{]}\\ &=\mathbb{E}\Big{[}\mathbb{E}\big{[}X_{N_{T}}\bm{1}_{X_{N_{T}}\leq r(N_{T})}\big{|}N_{T}\big{]}\Big{]}+\mathbb{E}\Big{[}\mathbb{E}\big{[}X_{N_{T}}\bm{1}_{X_{N_{T}}>r(N_{T})}\big{|}N_{T}\big{]}\Big{]}\\ &\leq\mathbb{E}\Big{[}\mathbb{E}\big{[}r(N_{T})\big{|}N_{T}\big{]}\mathbb{P}\big{(}X_{N_{T}}\leq r(N_{T})\big{|}N_{T}\big{)}\Big{]}\\ &\qquad+\mathbb{E}\Big{[}\mathbb{E}\big{[}X_{N_{T}}\big{|}N_{T},X_{N_{T}}>r(N_{T})\big{]}\mathbb{P}\big{(}X_{N_{T}}>r(N_{T})\big{|}N_{T}\big{)}\Big{]}\\ &\leq\mathbb{E}\big{[}C_{1}\ln^{2}N_{T}+C_{2}\ln N_{T}\ln(\tfrac{2}{\epsilon})+C_{3}\ln^{2}(\tfrac{2}{\epsilon})+C_{4}\big{]}\\ &\qquad+(\bar{K}-\underline{K})^{2}\mathbb{E}\Big{[}\sum_{n=0}^{N_{T}}(\tau_{n+1}-\tau_{n})(2\epsilon)\Big{]}\,.\end{split}$
Let us set $\epsilon=\frac{2}{T}$ and we can take $\epsilon$ out of the expectation. Besides, we know that $x\mapsto\ln x$ is concave and, for large $x$, i.e. $x\geq 3$, $x\mapsto\ln^{2}x2$ is concave, hence by Jensen’s inequality, we have
@@
where we use the fact that $\ln T\leq\ln^{2}T2$ for large $T$ and we ignore the term of order $\mathcal{O}(T^{-1})$. By using Theorem 27, we then have
$\begin{split}\mathcal{R}^{\hat{\Psi}}(T)&\leq C_{1}^{\prime}\mathbb{E}\Big{[}\int_{0}^{T}\left|\kappa_{t}-\kappa^{\star}\right|^{2}\,\mathrm{d}t\Big{]}+C_{2}^{\prime}\frac{e^{T\ln\alpha}-1}{\ln\alpha}\\ &\leq C_{1}^{\prime}\mathbb{E}\Big{[}X_{N_{T}}\Big{]}+\frac{C_{2}^{\prime}}{\ln\frac{1}{\alpha}}(1-\alpha^{T})\\ &\leq C_{1}^{\prime}\left(C_{1}+C_{2}(1+\ln(\lambda^{+}+\lambda^{-})+C_{3})\right)\ln^{2}T\\ &\qquad+C_{1}^{\prime}\left(C_{1}\ln^{2}(\lambda^{+}+\lambda^{-})+C_{4}+4(\bar{K}-\underline{K})^{2}\right)+\frac{C_{2}^{\prime}}{\ln\frac{1}{\alpha}}\,,\end{split}$
where $C_{1}^{\prime},C_{2}^{\prime}>0$ and $10<\alpha<$ 1 are constants independent of $T$ in Theorem  27 and $C_{1},C_{2},C_{3},C_{4}$ are constants independent of $\epsilon$ and $T$ in Proposition 34, hence the result of Theorem 19.

### v2

@@ 2.2. Learning and regret
There exist a constant $C>0$ depends on $\lambda^{\pm},\underline{K}$ and $\bar{K}$ such that
$0\leq\gamma(\kappa^{\ast};\kappa^{\ast})-\gamma(\kappa;\kappa^{\ast})\leq C\left|\kappa-\kappa^{\ast}\right|^{2},\,\forall\kappa\in[\underline{K},\bar{K}]\,.$
Remark 25.
The constant $C$ in Lemma 24 implicitly depends on the model parameters, such as $\lambda^{\pm},\underline{K}$ and $\bar{K}$. Although we prove that $\kappa\mapsto\gamma(\kappa;\kappa^{\ast})$ is twice continuously differentiable, we do not have an analytic expression as it depends on derivatives of the eigenvalues and eigenvectors of the matrix whose entries are functions of the model parameters. While this is no obstacle to asymptotic regret analysis it may be interesting to quantify the dependence of $C$ on the model parameters. This has been done numerically, see Figure 6.
So far we have performed the ergodic analysis for the market making model with the parameter $\kappa$ misspecified. Another key step towards quantifying the performance gap, see Theorem 30, is to analyse the ergodicity under the model misspecification, i.e. how fast the state process $\big{(}Q_{t}^{\psi^{\kappa};\kappa^{\ast}}\big{)}_{t\geq 0}$ following the dynamics (23) converges to the equilibrium distribution.
Definition 26 (Equilibrium).
The following lemma is proved in Appendix A.15.
Lemma 27.
For any $\kappa\in[\underline{K},\bar{K}]$, the controlled process $\big{(}Q_{t}^{\psi^{\kappa};\kappa^{\ast}}\big{)}_{t\geq 0}$, following the dynamics (23) under the control $\psi^{\kappa}$, admits a unique equilibrium distribution, denoted by $\pi^{\psi^{\kappa};\kappa^{\ast}}$.
As we show in Appendix A.15, $(Q_{t})_{t\geq 0}$–with superscripts omitted for brevity–can be equivalently represented as a continuous-time Markov chain (CTMC) with the transition rate matrix $\bm{Q}$ given by (60). Since the transition rate matrix $\bm{Q}$ is tridiagonal, the CTMC is irreducible and recurrent. Therefore, the convergence of the distribution of $Q_{t}$ to the equilibrium distribution follows the Convergence Theorem [28, Theorem 3.6].
Lemma 28 (Convergence Theorem).
Let $\pi_{t}^{\psi^{\kappa};\kappa^{\ast}}$ be the probability distribution of the random variables $Q_{t}$ that follows the controlled dynamics (23) with an initial state $Q_{0}\sim\pi_{0}$ and the control $\psi^{\kappa}$. Let $\pi^{\psi^{\kappa};\kappa^{\ast}}$ be the equilibrium distribution established by Lemma 27. Then, there exists constants $C>0$ and $10<\alpha<$ 1, depending on $\kappa$, such that
$\big{\|}\pi_{t}^{\psi^{\kappa};\kappa^{\ast}}-\pi^{\psi^{\kappa};\kappa^{\ast}}\big{\|}_{TV}\leq C\alpha^{t},\quad\forall t\geq 0\,.$
We next state the following proposition, proved in Appendix A.16, which establishes a key property of the equilibrium distribution $\pi^{\psi^{\kappa};\kappa^{\ast}}$.
@@
Let $\pi^{\psi^{\kappa};\kappa^{\ast}}$ be the equilibrium distribution for the inventory process $Q_{t}$ following the controlled SDE (23)
$dQ_{t}^{\psi^{\kappa};\kappa^{\ast}}=\big{(}\lambda^{+}e^{-\kappa^{\ast}\psi^{\kappa,-}}-\lambda^{-}e^{-\kappa^{\ast}\psi^{\kappa,+}}\big{)}dt+d\tilde{N}_{t}^{\psi^{\kappa,-}}-d\tilde{N}_{t}^{\psi^{\kappa,+}},\quad t\in[0,T],Q_{0}\sim\pi^{\psi^{\kappa};\kappa^{\ast}},$
@@
$\begin{split}&\mathbb{E}\bigg{[}\int_{0}^{T}\lambda^{+}e^{-\kappa^{\ast}\psi^{\kappa,+}(Q_{t}^{\psi^{\kappa};\kappa^{\ast}})}\Big{(}\hat{v}^{\psi^{\kappa}}(Q_{t}^{\psi^{\kappa};\kappa^{\ast}}-1;\kappa^{\ast})-\hat{v}^{\psi^{\kappa}}(Q_{t}^{\psi^{\kappa};\kappa^{\ast}};\kappa^{\ast})\Big{)}\mathds{1}_{Q_{t}^{\psi^{\kappa};\kappa^{\ast}}>\underline{q}}\\ &+\lambda^{-}e^{-\kappa^{\ast}\psi^{\kappa,-}(Q_{t}^{\psi^{\kappa};\kappa^{\ast}})}\Big{(}\hat{v}^{\psi^{\kappa}}(Q_{t}^{\psi^{\kappa};\kappa^{\ast}}+1;\kappa^{\ast})-\hat{v}^{\psi^{\kappa}}(Q_{t}^{\psi^{\kappa};\kappa^{\ast}};\kappa^{\ast})\Big{)}\mathds{1}_{Q_{t}^{\psi^{\kappa};\kappa^{\ast}}<\bar{q}}\,\mathrm{d}t\bigg{]}=0\,,\end{split}$
where $\hat{v}^{\psi^{\kappa}}(q;\kappa^{\ast})$ is defined in Proposition 19.
@@
$\mathcal{R}^{\Psi}(T)\leq C_{1}\mathbb{E}\Big{[}\int_{0}^{T}\left|\kappa_{t}-\kappa^{\ast}\right|^{2}\,\mathrm{d}t\Big{]}+\frac{C_{2}}{\ln(\alpha^{-1})}\,,$
@@
The next step towards Theorem 17 is to quantify the precise tail behaviour, also known as concentration inequality, of the regularised maximum-likelihood estimator in Algorithm 1. Recall that $(\Omega^{\ast},\mathcal{F}^{\ast},\mathbb{P}^{\ast})$ is given in Definition 11.
We start with several significant propositions to the estimator. The proofs of the following propositions are provided in Appendix A.18.
@@
Let $(\delta_{n})_{n=1}^{N}$ be a collection of non-negative random variables taking values in $[\underline{\delta},\bar{\delta}]\cup\{+\infty\}$. Then for any $\varepsilon\geq 0$ and bounded function $f:[\underline{\delta},\bar{\delta}]\cup\{+\infty\}\rightarrow\mathbb{R}$, it holds that,
$\mathbb{P}^{\ast}\left(\left|\sum_{n=1}^{N}f(\delta_{n})Y_{n}-\sum_{n=1}^{N}f(\delta_{n})e^{-\kappa^{\ast}\delta_{n}}\right|\leq\left\|f\right\|_{\infty}\sqrt{2N\ln(\tfrac{2}{\varepsilon})}\right)\geq 1-\varepsilon.$
@@
There exist constants $c,C>0$ depending on $\underline{K},\bar{K},\underline{\delta}$ and $\bar{\delta}$ such that for any policy $(\delta_{n})_{n=1}^{\infty}$ taking values in $[\underline{\delta},\bar{\delta}]\cup\{+\infty\}$, it holds that for any $\varepsilon>0$,
$\mathbb{P}^{\ast}\left(\inf_{\kappa>0}\Big{(}-\frac{\mathrm{d}^{2}}{\mathrm{d}\kappa^{2}}\tilde{\ell}_{N}(\kappa)\Big{)}\geq cN-C\sqrt{N\ln\big{(}\tfrac{2}{\varepsilon}\big{)}}\right)\geq 1-\varepsilon.$
@@
There exists constants $C,c\geq 0$ depending on $\kappa^{\ast}$, $\delta_{0}$, and $\bar{\delta}$ such that for any policy $(\delta_{n})_{n=1}^{N}$ taking values in $[\underline{\delta},\bar{\delta}]\cup\{+\infty\}$, it holds that for any $\varepsilon>0$,
$\mathbb{P}^{\ast}\left(\Big{|}\frac{\mathrm{d}}{\mathrm{d}\kappa}\tilde{\ell}_{N}(\kappa^{\ast})\Big{|}\leq C\sqrt{N\ln\big{(}\tfrac{2}{\varepsilon}\big{)}}+c\right)\geq 1-\varepsilon.$
@@
Let $\kappa_{N}>0$ be the unique solution to $\frac{\mathrm{d}}{\mathrm{d}\kappa}\tilde{\ell}_{N}(\kappa_{N})=0$. There exists constants $C,c,N_{0}\geq 0$ such that for any $\varepsilon\geq 0$, if $N\geq N_{0}\ln\big{(}\tfrac{2}{\varepsilon}\big{)}$, then
$\mathbb{P}^{\ast}\left(|\kappa_{N}-\kappa^{\ast}|\leq CN^{-1/2}\sqrt{\ln\big{(}\tfrac{2}{\varepsilon}\big{)}}+cN^{-1}\right)\geq 1-2\varepsilon.$
@@
Let $\kappa_{N}>0$ be the unique solution to $\frac{\mathrm{d}}{\mathrm{d}\kappa}\tilde{\ell}_{N}(\kappa_{N})=0$. Then there exists constants $C,c,N_{0}\geq 0$ such that for any $\varepsilon\geq 0$,
$\mathbb{P}^{\ast}\left(|\kappa_{N}-\kappa^{\ast}|\leq CN^{-1/2}\sqrt{\ln\big{(}\tfrac{2N}{\varepsilon}\big{)}}+cN^{-1}\quad\text{for all}\quad N\geq N_{0}\ln\big{(}\tfrac{2}{\varepsilon}\big{)}\right)\geq 1-\varepsilon\,.$
Another corollary to the above results, which implies that for sufficiently large $N$ the estimator will eventually remain within the compact set $[\underline{K},\bar{K}]$, is stated below. See Appendix A.20 for the proof.
Corollary 35.2.
Let $\kappa_{N}>0$ be the unique solution to $\frac{\mathrm{d}}{\mathrm{d}\kappa}\tilde{\ell}_{N}(\kappa_{N})=0$. Then there exists constants $N_{0},N_{0}^{\prime}\geq 0$ such that for any $\varepsilon\geq 0$
$\mathbb{P}^{\ast}\left(\kappa_{N}\in[\underline{K},\bar{K}]\quad\text{for all}\quad N\geq\max\left(N_{0}\ln\big{(}\tfrac{2}{\varepsilon}\big{)},N_{0}^{\prime}/\ln\big{(}\tfrac{2}{\varepsilon}\big{)}\right)\right)\geq 1-\varepsilon\,.$
@@ 2.2.1. Step 3: Proof of Theorem 17
$\int_{0}^{T}|\kappa_{t}-\kappa^{\ast}|^{2}\,\mathrm{d}t\leq\sum_{n=0}^{N_{T}}(\tau_{n+1}-\tau_{n})|\kappa_{n}-\kappa^{\ast}|^{2}=:X_{N_{T}}\,.$
@@
The following proposition, Proposition 36, which is proved in Appendix  A.21, states that, given any $N_{T}$, i.e. the number of signals up to time $T$, the random variable $X_{N_{T}}$ is bounded by $\mathcal{O}(\ln^{2}N_{T})$ with high probability.
@@
There exist constants $C_{1},C_{2},C_{3},C_{4}>0$ such that for any $\varepsilon>0$,
$\mathbb{P}^{\ast}\left(X_{N_{T}}\leq C_{1}\ln^{2}N_{T}+C_{2}\ln N_{T}\ln(\tfrac{2}{\varepsilon})+C_{3}\ln^{2}(\tfrac{2}{\varepsilon})+C_{4}\right)\geq 1-2\varepsilon.$
Let $r(N_{T}):=C_{1}\ln^{2}N_{T}+C_{2}\ln N_{T}\ln(\tfrac{2}{\varepsilon})+C_{3}\ln^{2}(\tfrac{2}{\varepsilon})+C_{4}$. Then, by using Proposition 36, we have
$\begin{split}\mathbb{E}[X_{N_{T}}]&=\mathbb{E}\Big{[}\mathbb{E}\big{[}X_{N_{T}}\big{|}N_{T}\big{]}\Big{]}\\ &=\mathbb{E}\Big{[}\mathbb{E}\big{[}X_{N_{T}}\mathds{1}_{X_{N_{T}}\leq r(N_{T})}\big{|}N_{T}\big{]}\Big{]}+\mathbb{E}\Big{[}\mathbb{E}\big{[}X_{N_{T}}\mathds{1}_{X_{N_{T}}>r(N_{T})}\big{|}N_{T}\big{]}\Big{]}\\ &\leq\mathbb{E}\Big{[}\mathbb{E}\big{[}r(N_{T})\big{|}N_{T}\big{]}\mathbb{P}\big{(}X_{N_{T}}\leq r(N_{T})\big{|}N_{T}\big{)}\Big{]}\\ &\qquad+\mathbb{E}\Big{[}\mathbb{E}\big{[}X_{N_{T}}\big{|}N_{T},X_{N_{T}}>r(N_{T})\big{]}\mathbb{P}\big{(}X_{N_{T}}>r(N_{T})\big{|}N_{T}\big{)}\Big{]}\\ &\leq\mathbb{E}\big{[}C_{1}\ln^{2}N_{T}+C_{2}\ln N_{T}\ln(\tfrac{2}{\varepsilon})+C_{3}\ln^{2}(\tfrac{2}{\varepsilon})+C_{4}\big{]}\\ &\qquad+(\bar{K}-\underline{K})^{2}\mathbb{E}\Big{[}\sum_{n=0}^{N_{T}}(\tau_{n+1}-\tau_{n})(2\varepsilon)\Big{]}\,.\end{split}$
Let us set $\varepsilon=\frac{2}{T}$ and we can take $\varepsilon$ out of the expectation. Besides, we know that $x\mapsto\ln x$ is concave and, for large $x$, i.e. $x\geq 3$, $x\mapsto\ln^{2}x$ is concave, hence by Jensen’s inequality, we have
@@
where we use the fact that $\ln T\leq\ln^{2}T$ for large $T$ and we ignore the term of order $\mathcal{O}(T^{-1})$. By using Theorem 30, we then have
$\begin{split}\mathcal{R}^{\hat{\Psi}}(T)&\leq C_{1}^{\prime}\mathbb{E}\Big{[}\int_{0}^{T}\left|\kappa_{t}-\kappa^{\ast}\right|^{2}\,\mathrm{d}t\Big{]}+C_{2}^{\prime}\frac{e^{T\ln\alpha}-1}{\ln\alpha}\\ &\leq C_{1}^{\prime}\mathbb{E}\Big{[}X_{N_{T}}\Big{]}+\frac{C_{2}^{\prime}}{\ln(\alpha^{-1})}(1-\alpha^{T})\\ &\leq C_{1}^{\prime}\left(C_{1}+C_{2}(1+\ln(\lambda^{+}+\lambda^{-})+C_{3})\right)\ln^{2}T\\ &\qquad+C_{1}^{\prime}\left(C_{1}\ln^{2}(\lambda^{+}+\lambda^{-})+C_{4}+4(\bar{K}-\underline{K})^{2}\right)+\frac{C_{2}^{\prime}}{\ln(\alpha^{-1})}\,,\end{split}$
where $C_{1}^{\prime},C_{2}^{\prime}>0$ and $10<\alpha<$ 1 are constants independent of $T$ in Theorem  30 and $C_{1},C_{2},C_{3},C_{4}$ are constants independent of $\varepsilon$ and $T$ in Proposition 36, hence the result of Theorem 17.